OpenStudy (anonymous):

Solve the elimination method.(my number three key went out) 5a+2b= -three9 -5a+b=48 the solution is?

OpenStudy (eamier):

add equation 1 and equation 2. a will be eliminated

OpenStudy (eamier):

you got 3b=9

OpenStudy (anonymous):

so it would be three b=9

OpenStudy (anonymous):

ok so b=three

OpenStudy (eamier):

good

OpenStudy (anonymous):

so the 2nd part i have to solve for a

OpenStudy (anonymous):

to get 2 solution (9,?)

OpenStudy (eamier):

yes. substitute b=3 in any equation. prefer easy one

OpenStudy (eamier):

take second eq. -5a+3=48 -5a=45 a=-9

OpenStudy (anonymous):

so (9,-9)

OpenStudy (anonymous):

so this solution is infintly many solution

OpenStudy (anonymous):

i see how you got the 48 divide into 5=-9

OpenStudy (eamier):

only one answer for a and b. a=-9,b=3

OpenStudy (anonymous):

oh ok we got the first answer three ok ok

OpenStudy (anonymous):

thanks